OpenStudy (anonymous):

Find all real solution: sqrt of x+2= x+5

OpenStudy (precal):

\[\sqrt{x+2}=x+5\] take both sides to the 2 power

OpenStudy (anonymous):

If u square both sides u get x+2=x^2+10x+25 x^2+9x+23=0 U wont get any real solution.

OpenStudy (anonymous):

Okay that's what I was stuck at becaue I know that doesn't factor I just didn't know if I had to put no real solution but now I do! Thanks a lot

OpenStudy (anonymous):

yw